2. An airplane travels 500 m down the runway before taking off. If it starts from rest, moves with constant acceleration, and becomes airborne in 20 s, with what velocity does it take off?
1
Expert's answer
2020-12-06T17:19:28-0500
"v=at=\frac{2s}{t^2}t=\frac{2s}{t}\\v=\frac{2(500)}{20}=50\frac{m}{s}"